OpenStudy (anonymous):

solve the following system of the equations x+4y=3 (1) x=9-4y9 (2)

OpenStudy (saifoo.khan):

what's -4y9?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

X+4y=3 (1) x=9-4y (2)

OpenStudy (saifoo.khan):

no solution. the lines wont intersect at any point.